# SmartBugBert: BERT-Enhanced Vulnerability Detection for Smart Contract Bytecode

### v1

@@ 1. Introduction
With the continuous expansion of application scenarios, the number of smart contracts deployed on the blockchain shows explosive growth (Zheng et al., 2023). Due to the irreversibility of blockchain, it is difficult to repair the vulnerabilities of deployed smart contracts (Chen et al., 2020). This makes the security of on-chain smart contracts face serious challenges (Xia et al., 2021). In order to verify the correctness of smart contracts and reduce the losses caused by security issues, a method that can efficiently detect smart contract vulnerabilities is essential (Luu et al., 2016).
@@
As shown in Fig. 1, some smart contract vulnerability detection efforts are implemented based on source code (Feist et al., 2019). While smart contracts deployed on blockchain systems are publicly transparent, it is not mandatory for contract developers to publish the source code. As a result, less than four of smart contracts on Ether are open source (Garamvölgyi et al., 2022). Although some studies have been conducted to implement smart contract vulnerability detection from a bytecode perspective, a simple piece of bytecode or opcode is difficult to provide explicit vulnerability characterization, limiting its effectiveness in smart contract vulnerability detection (Tsankov et al., 2018). To overcome this limitation, this paper extracts CFG from smart contract bytecode, which contains rich smart contract business logic and thus helps to realize BERT-based smart contract detection work more effectively.
In order to be able to more effectively detect vulnerabilities in bytecode-level smart contracts, this paper designs SmartBugBert, an efficient smart contract vulnerability detection method based on the BERT extension. To further improve the detection effect of smart contract vulnerabilities, this paper also combines the control flow graph in the static analysis technique and integrates the multi-dimensional detection method. The method is capable of detecting vulnerabilities from the smart contract bytecode and identifies four vulnerabilities: reentry vulnerability, arithmetic vulnerability, self-destructing contract, and timestamp dependency vulnerability (Zhuang et al., 2020).
@@ 2.1. Smart Contract Vulnerabilities
Smart contracts are self-executing digital agreements written in code that automatically enforce and execute predefined terms when specific conditions are met (Wood et al., 2014). These contracts operate on blockchain platforms and facilitate decentralized applications (DApps) and decentralized finance (DeFi) protocols (Li et al., 2024d, c). However, due to their immutable nature, smart contract vulnerabilities can lead to significant financial losses if exploited (Li et al., 2024b; Zheng et al., 2023).
@@
  • Transaction-Ordering Vulnerability (TOV): This vulnerability arises when the execution result of a transaction depends on the order in which transactions are mined, allowing attackers to manipulate transaction execution sequences for profit (Li et al., 2024a; Zhuang et al., 2020).
  • Access Control Vulnerability (ACV): This occurs when sensitive contract functions lack proper authorization checks, potentially allowing unauthorized users to execute privileged operations (Li et al., 2024b; Ghaleb et al., 2023).
  • Self-Destruct Vulnerability (SDV): This vulnerability enables attackers to trigger a contract’s self-destruct mechanism inappropriately, which can lead to permanent deletion of the contract and its assets (Li et al., 2020; Tikhomirov et al., 2018).
  • Timestamp Dependency Vulnerability (TDV): This vulnerability exists when contracts rely on block timestamps for critical operations, which miners can manipulate within certain bounds (Li et al., 2024a; Nikolić et al., 2018).
The prevalence of these vulnerabilities in smart contracts, especially in emerging ecosystems like NFT marketplaces (Kong et al., 2024; Niu et al., 2024), highlights the urgent need for effective detection techniques (Yang et al., 2023; Chen et al., 2020).
@@ 2.2. Control Flow Graphs for Smart Contracts
A CFG is a representation of all paths that might be traversed through a program during its execution (Mossberg et al., 2019). In the context of smart contracts, CFGs provide valuable structural information about the contract’s execution flow and help identify potential vulnerability patterns (Li et al., 2024a).
Traditional approaches to vulnerability detection often rely solely on bytecode information without considering the control flow structure, which can lead to false positives or missed vulnerabilities (Feist et al., 2019). Our approach addresses this limitation by recovering the CFG from the contract bytecode and extracting vulnerable CFG fragments that contain potential vulnerabilities. This approach enables more targeted and efficient vulnerability detection compared to analyzing the entire contract code (Mao et al., 2024; Xu et al., 2023).
@@ 2.3. BERT-based Smart Contract Analysis
Recent advancements in natural language processing, particularly the Bidirectional Encoder Representations from Transformers (BERT) model, have shown promising results in code analysis tasks (Li et al., 2023; Wang et al., 2022). BERT’s ability to capture contextual relationships in sequences makes it well-suited for analyzing program code and identifying complex patterns (Asudani et al., 2023).
For smart contract vulnerability detection, BERT can be fine-tuned to extract features from the CFG that represent potential vulnerable patterns (Li et al., 2024a; Samreen and Alalfi, 2021). This approach offers advantages over traditional feature engineering methods as it can automatically learn relevant features from the data (Tann et al., 2018).
Our work builds upon these foundations by combining BERT-based feature extraction with statistical semantic features and using LightGBM for classification. This integrated approach allows for more comprehensive vulnerability detection that considers both the semantic context and the control flow structure of smart contracts (Durieux et al., 2020; di Angelo et al., 2023).
@@ 3. Method
In this section, SmartBugBERT smart contract vulnerability detection method is proposed, which mainly consists of three major parts: semantic extraction module, bytecode-level CFG module construction, and CFG vulnerability fragment extraction module. Specifically, the implementation steps of SmartBugBERT are shown in Fig. 2: The context information, i.e., opcode information, is extracted from the collected bytecode-level smart contracts using (1) decompilation module (Li, 2021). After filtering the opcodes with the same function in the opcode information and (2) extracting the semantics; from the sequence of opcodes obtained from the decompiler module, (3) construct the control flow graph of the smart contract and extract the CFG fragments with vulnerabilities through (4) vulnerability fragments, utilize the Bert feature extraction and fuse it with the semantic features; the fused full features are sent to the classification module to complete the (5) vulnerability detection task and generate the report.
@@ 3.1. Semantic Extraction Module
Bytecode is stored on the blockchain as a string of hexadecimal numbers. Unlike source code, bytecode is completely open and transparent and can be easily accessed from each contract (Li et al., 2017). The bytecode-based semantic extraction module has two major steps: bytecode decompilation and feature extraction.
First, consider that smart contract bytecode is not easy to read for humans and does not have any semantic information. In this paper, we convert the bytecode into equivalent opcodes that are easy to be understood by humans through the pyevmasm disassembler to facilitate semantic extraction of the contract (Li et al., 2024e). As shown in Fig. 3, the initial sequence of opcodes is redundant, and for better semantic extraction, SmartBugBERT optimizes the representation of operands with the same behavior, e.g., DUP1 and DUP2 are both considered as DUP; PUSH1 and PUSH2 are both considered as PUSH (Li et al., 2025).
Second, the optimized opcodes are statistically measured using TF-IDF, which describes the importance of a given opcode in a certain vulnerability category.TF-IDF measures the importance of a single opcode by the product of two parameters, the word frequency (TF) and the inverse document frequency (IDF).TF reflects the frequency of occurrence of a word in a document, while the IDF describes the rarity of a single opcode in the entire rarity of a single operand in the entire document collection. In general, the closer the IDF value is to 0, the more common the word is, and conversely, the more representative the word corresponding to the operand is (Wang et al., 2024).
For example, among them PUSH, DUP, SWAP and POP are the four most commonly used in smart contracts. These opcodes are all related to stack operations. Since EVMs are stack-based, almost any operation, such as defining variables and functions, performing arithmetic operations (pressing data into the EVM stack), swapping elements, and deleting variables, requires stack operations. As a result, the IDFs of these opcodes are almost semantically unimportant. However, the IDF value of the opcode SELFDESTRUCT (byte value 0xFF) in a contract with a self-destruct vulnerability will exceed the IDF value of a contract with other vulnerabilities (Liu et al., 2024). Therefore, the statistical characterization of opcodes can be used to detect vulnerabilities in contracts, and it reflects the characterization of contract vulnerabilities to some extent from an EVM perspective.
@@ 3.2. Control Flow Diagram Building Blocks at the Byte-Code Level
17:         // Set the previous instruction as the end of bb
18:         Set end instruction of bb to previous op
19:         Append bb to basic_block_list
@@
25:         // Set current instruction as the end of bb
26:         Set end instruction of bb to op
27:         Append bb to basic_block_list
@@
34:for each bb in basic_block_list do
35:     AddEdge(bb)
@@ 4.1. Experimental Data
Since current research does not provide annotated datasets for bytecode-level smart contract vulnerability detection, and to ensure the authenticity of SmartBugBERT’s effectiveness, real smart contracts were collected from Ethersca and annotated using existing smart contract detection tools. First, Google BigQuery was used to collect 14,289 smart contract addresses. Then, Python scripts were employed to request the Solidity source code files and bytecode files of these smart contracts from Etherscan, establishing corresponding relationships. After filtering out duplicate contracts and those without source code, 9,346 unique bytecode contract files were obtained. Finally, existing contract vulnerability detection tools were used to annotate these contracts. Due to differences in the capabilities of various tools, we selected more advanced tools to collect contracts with RA, AV, SD, and TDV vulnerabilities as accurately as possible. Oyente was used to identify RA and TDV vulnerabilities, MAIAN to identify SD vulnerabilities, and Osiris to identify AV vulnerabilities. After filtering out contracts without detected vulnerabilities, 6,157 processed contracts and their corresponding vulnerability labels were collected.
@@ 4.4. Experimental Results and Analysis
To further evaluate the contract vulnerability detection effectiveness of our method, we compared it with two other smart contract vulnerability detection approaches: SaferSC and Oyente. As shown in Table 1, our method achieved significant improvements in vulnerability detection compared to both SaferSC and Oyente. Specifically, our method improved precision by 41.19% over SaferSC and 48.86% over Oyente, while improving recall by 39.51% over SaferSC and 45.25% over Oyente. The relatively poor performance of Oyente is attributed to its inability to detect self-destructing contracts, which reflects the limitations of traditional contract vulnerability detection methods.
@@ 4.5. Ablation Study
Table 4 shows that the average vulnerability detection time for the symbolic execution tool is 528.57 seconds, while the proposed machine learning-based SaferSC and SmartBugBert methods have average detection times of 0.14 seconds and 0.23 seconds respectively. This difference arises because machine learning methods only detect specific vulnerabilities they were trained on, without performing comprehensive analysis of other characteristics of smart contracts.
@@ 5. Conclusion
This paper introduces a BERT-based bytecode-level smart contract vulnerability detection method. The approach first decompiles smart contract bytecode into opcode sequences and represents their semantic features through statistical characteristics. Unlike methods that solely use bytecode information, this chapter recovers the CFG from the bytecode level and extracts vulnerable CFG fragments containing potential vulnerabilities for efficient detection, while utilizing fine-tuned BERT to extract CFG features. Subsequently, the semantic features and CFG features are fused as input to the LightGBM classifier to accomplish the contract vulnerability detection task. Experimental results demonstrate that the proposed method can effectively detect Transaction-Ordering Vulnerability, Access Control Vulnerability, Self-Destruct Vulnerability, and Timestamp Dependency Vulnerability in contracts, achieving an excellent F1-Score of 91.19%. Regarding detection time, it also shows significant advantages compared to symbolic execution-based tools.

### v2

@@ 1. Introduction
With the continuous expansion of application scenarios, the number of smart contracts deployed on the blockchain shows explosive growth (Zheng_2023_Blockchain-Based). Due to the irreversibility of blockchain, it is difficult to repair the vulnerabilities of deployed smart contracts (Chen_2020_Survey). This makes the security of on-chain smart contracts face serious challenges (Xia_2021_Tradea). In order to verify the correctness of smart contracts and reduce the losses caused by security issues, a method that can efficiently detect smart contract vulnerabilities is essential (Luu_2016_Making).
@@
As shown in Fig. 1, some smart contract vulnerability detection efforts are implemented based on source code (Feist_2019_Slither; brent2020ethainter; schneidewind2020ethor). While smart contracts deployed on blockchain systems are publicly transparent, it is not mandatory for contract developers to publish the source code. As a result, less than four of smart contracts on Ether are open source (Garamvolgyi_2022_Utilizing). Although some studies have been conducted to implement smart contract vulnerability detection from a bytecode perspective (brent2018vandal; Veloso__Conkas; 10.1145/3543507.3583367), a simple piece of bytecode or opcode is difficult to provide explicit vulnerability characterization, limiting its effectiveness in smart contract vulnerability detection (tsankov2018securify). To overcome this limitation, this paper extracts CFG from smart contract bytecode, which contains rich smart contract business logic and thus helps to realize BERT-based smart contract detection work more effectively.
In order to be able to more effectively detect vulnerabilities in bytecode-level smart contracts, this paper designs SmartBugBert, an efficient smart contract vulnerability detection method based on the BERT extension. To further improve the detection effect of smart contract vulnerabilities, this paper also combines the control flow graph in the static analysis technique and integrates the multi-dimensional detection method. The method is capable of detecting vulnerabilities from the smart contract bytecode and identifies four vulnerabilities: reentry vulnerability, arithmetic vulnerability, self-destructing contract, and timestamp dependency vulnerability (Zhuang_2020_Smart; luo2024scvhunter; zhen2024gnn).
@@ 2.1. Smart Contract Vulnerabilities
Smart contracts are self-executing digital agreements written in code that automatically enforce and execute predefined terms when specific conditions are met (Ethereum). These contracts operate on blockchain platforms and facilitate decentralized applications (DApps) and decentralized finance (DeFi) protocols (li2024defitail; li2024stateguard). However, due to their immutable nature, smart contract vulnerabilities can lead to significant financial losses if exploited (li2024guardians; Zheng_2023_Blockchain-Based; slowmist).
@@
  • Transaction-Ordering Vulnerability (TOV): This vulnerability arises when the execution result of a transaction depends on the order in which transactions are mined, allowing attackers to manipulate transaction execution sequences for profit (li2024cobra; Zhuang_2020_Smart).
  • Access Control Vulnerability (ACV): This occurs when sensitive contract functions lack proper authorization checks, potentially allowing unauthorized users to execute privileged operations (li2024guardians; Ghaleb_2023_AChecker).
  • Self-Destruct Vulnerability (SDV): This vulnerability enables attackers to trigger a contract’s self-destruct mechanism inappropriately, which can lead to permanent deletion of the contract and its assets (li2020characterizing; Tikhomirov_2018_SmartCheck; torres2019art).
  • Timestamp Dependency Vulnerability (TDV): This vulnerability exists when contracts rely on block timestamps for critical operations, which miners can manipulate within certain bounds (li2024cobra; Nikolic_2018_Finding; Luu_2016_Making).
The prevalence of these vulnerabilities in smart contracts, especially in emerging ecosystems like NFT marketplaces (kong2024characterizing; niu2024unveiling; ma2025uncovering), highlights the urgent need for effective detection techniques (Yang_2023_Definition; Chen_2020_Survey).
@@ 2.2. Control Flow Graphs for Smart Contracts
A CFG is a representation of all paths that might be traversed through a program during its execution (Mossberg_2019_Manticore; wang2023smart). In the context of smart contracts, CFGs provide valuable structural information about the contract’s execution flow and help identify potential vulnerability patterns (li2024cobra).
Traditional approaches to vulnerability detection often rely solely on bytecode information without considering the control flow structure, which can lead to false positives or missed vulnerabilities (Feist_2019_Slither; DurieuxEtAl2020ICSE). Our approach addresses this limitation by recovering the CFG from the contract bytecode and extracting vulnerable CFG fragments that contain potential vulnerabilities. This approach enables more targeted and efficient vulnerability detection compared to analyzing the entire contract code (mao2024scla; Xu_2023_SoK).
@@ 2.3. BERT-based Smart Contract Analysis
Recent advancements in natural language processing, particularly the Bidirectional Encoder Representations from Transformers (BERT) model, have shown promising results in code analysis tasks (li2023overview; Wang_2022_Unified). BERT’s ability to capture contextual relationships in sequences makes it well-suited for analyzing program code and identifying complex patterns (Asudani_2023_Impact).
For smart contract vulnerability detection, BERT can be fine-tuned to extract features from the CFG that represent potential vulnerable patterns (li2024cobra; Samreen_2021_SmartScan; 10646885). This approach offers advantages over traditional feature engineering methods as it can automatically learn relevant features from the data (tann2019safer).
Our work builds upon these foundations by combining BERT-based feature extraction with statistical semantic features and using LightGBM for classification. This integrated approach allows for more comprehensive vulnerability detection that considers both the semantic context and the control flow structure of smart contracts (DurieuxEtAl2020ICSE; diAngeloEtAl2023ASE; chen2024improving; shang2025cegt; 10700860).
@@ 3. Method
In this section, SmartBugBERT smart contract vulnerability detection method is proposed, which mainly consists of three major parts: semantic extraction module, bytecode-level CFG module construction, and CFG vulnerability fragment extraction module. Specifically, the implementation steps of SmartBugBERT are shown in Fig. 2: The context information, i.e., opcode information, is extracted from the collected bytecode-level smart contracts using (1) decompilation module (li2021hybrid). After filtering the opcodes with the same function in the opcode information and (2) extracting the semantics; from the sequence of opcodes obtained from the decompiler module, (3) construct the control flow graph of the smart contract and extract the CFG fragments with vulnerabilities through (4) vulnerability fragments, utilize the Bert feature extraction and fuse it with the semantic features; the fused full features are sent to the classification module to complete the (5) vulnerability detection task and generate the report.
@@ 3.1. Semantic Extraction Module
Bytecode is stored on the blockchain as a string of hexadecimal numbers. Unlike source code, bytecode is completely open and transparent and can be easily accessed from each contract (li2017discovering). The bytecode-based semantic extraction module has two major steps: bytecode decompilation and feature extraction.
First, consider that smart contract bytecode is not easy to read for humans and does not have any semantic information. In this paper, we convert the bytecode into equivalent opcodes that are easy to be understood by humans through the pyevmasm disassembler to facilitate semantic extraction of the contract (li2024detecting). As shown in Fig. 3, the initial sequence of opcodes is redundant, and for better semantic extraction, SmartBugBERT optimizes the representation of operands with the same behavior, e.g., DUP1 and DUP2 are both considered as DUP; PUSH1 and PUSH2 are both considered as PUSH (li2025scalm).
Second, the optimized opcodes are statistically measured using TF-IDF, which describes the importance of a given opcode in a certain vulnerability category.TF-IDF measures the importance of a single opcode by the product of two parameters, the word frequency (TF) and the inverse document frequency (IDF).TF reflects the frequency of occurrence of a word in a document, while the IDF describes the rarity of a single opcode in the entire rarity of a single operand in the entire document collection. In general, the closer the IDF value is to 0, the more common the word is, and conversely, the more representative the word corresponding to the operand is (wang2024smart).
For example, among them PUSH, DUP, SWAP and POP are the four most commonly used in smart contracts. These opcodes are all related to stack operations. Since EVMs are stack-based, almost any operation, such as defining variables and functions, performing arithmetic operations (pressing data into the EVM stack), swapping elements, and deleting variables, requires stack operations. As a result, the IDFs of these opcodes are almost semantically unimportant. However, the IDF value of the opcode SELFDESTRUCT (byte value 0xFF) in a contract with a self-destruct vulnerability will exceed the IDF value of a contract with other vulnerabilities (10707457). Therefore, the statistical characterization of opcodes can be used to detect vulnerabilities in contracts, and it reflects the characterization of contract vulnerabilities to some extent from an EVM perspective.
@@ 3.2. Control Flow Diagram Building Blocks at the Byte-Code Level
17:   // Set the previous instruction as the end of $b$
18:   Set end instruction of $b$ to previous op
19:   Append $b$ to basic_block_list
@@
25:   // Set current instruction as the end of $b$
26:   Set end instruction of $b$ to op
27:   Append $b$ to basic_block_list
@@
34:for each $b$ in basic_block_list do
35:  AddEdge($b$)
@@ 4.1. Experimental Data
Since current research does not provide annotated datasets for bytecode-level smart contract vulnerability detection, and to ensure the authenticity of SmartBugBERT’s effectiveness, real smart contracts were collected from Ethersca and annotated using existing smart contract detection tools (10486822; LI2026122645). First, Google BigQuery was used to collect 14,289 smart contract addresses. Then, Python scripts were employed to request the Solidity source code files and bytecode files of these smart contracts from Etherscan, establishing corresponding relationships. After filtering out duplicate contracts and those without source code, 9,346 unique bytecode contract files were obtained. Finally, existing contract vulnerability detection tools were used to annotate these contracts. Due to differences in the capabilities of various tools, we selected more advanced tools to collect contracts with RA, AV, SD, and TDV vulnerabilities as accurately as possible. Oyente was used to identify RA and TDV vulnerabilities (Luu_2016_Making), MAIAN to identify SD vulnerabilities, and Osiris to identify AV vulnerabilities (torres2018osiris). After filtering out contracts without detected vulnerabilities, 6,157 processed contracts and their corresponding vulnerability labels were collected.
@@ 4.4. Experimental Results and Analysis
To further evaluate the contract vulnerability detection effectiveness of our method, we compared it with two other smart contract vulnerability detection approaches: SaferSC and Oyente (Luu_2016_Making; grieco2020echidna; wustholz2020harvey). As shown in Table 1, our method achieved significant improvements in vulnerability detection compared to both SaferSC and Oyente. Specifically, our method improved precision by 41.19% over SaferSC and 48.86% over Oyente, while improving recall by 39.51% over SaferSC and 45.25% over Oyente. The relatively poor performance of Oyente is attributed to its inability to detect self-destructing contracts, which reflects the limitations of traditional contract vulnerability detection methods.
@@ 4.5. Ablation Study
Table 4 shows that the average vulnerability detection time for the symbolic execution tool is 528.57 seconds, while the proposed machine learning-based SaferSC and SmartBugBert methods have average detection times of 0.14 seconds and 0.23 seconds respectively. This difference arises because machine learning methods only detect specific vulnerabilities they were trained on, without performing comprehensive analysis of other characteristics of smart contracts (mythril; Mossberg_2019_Manticore; chen2025chatgpt; mueller2018smashing; pakala).
@@ 5. Conclusion
This paper introduces a BERT-based bytecode-level smart contract vulnerability detection method (10.1145/3699597; song2025silence). The approach first decompiles smart contract bytecode into opcode sequences and represents their semantic features through statistical characteristics. Unlike methods that solely use bytecode information, this chapter recovers the CFG from the bytecode level and extracts vulnerable CFG fragments containing potential vulnerabilities for efficient detection, while utilizing fine-tuned BERT to extract CFG features. Subsequently, the semantic features and CFG features are fused as input to the LightGBM classifier to accomplish the contract vulnerability detection task. Experimental results demonstrate that the proposed method can effectively detect Transaction-Ordering Vulnerability, Access Control Vulnerability, Self-Destruct Vulnerability, and Timestamp Dependency Vulnerability in contracts, achieving an excellent F1-Score of 91.19% (liu2025detecting; RuntimeVerification). Regarding detection time, it also shows significant advantages compared to symbolic execution-based tools.